OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4?

OpenStudy (anonymous):

y=mx+c

OpenStudy (anonymous):

m=4 y=4x + c

OpenStudy (anonymous):

This are the options : a. y - 1 = 4(x - 3) b. y + 1 = 4(x + 3) c. y - 3 = 4(x + 1) d. y + 3 = 4(x - 1)

OpenStudy (anonymous):

to find c substitute y and x -1= 4*-3 + c

OpenStudy (anonymous):

c= 11 y=4x +11

OpenStudy (anonymous):

y+1 = 4(x+3)

OpenStudy (anonymous):

So the answer is 4(x+3)?

OpenStudy (anonymous):

Okay i got it thanks .

OpenStudy (anonymous):

Can somebody help with another question please ?